# Weak Fano threefolds arising as the blowup of a hyperquadric in $\mathbb{P}^{4}$ along a curve

### v1

@@ 4. Curves on a hyperplane or smooth hyperquadric section
Note that there are $|\mathcal{P}|=36$ distinct $(g,d)$-pairs corresponding to $|\mathcal{P}_{\textup{plane}}|+|\mathcal{P}_{\textup{quadric}}|+|\mathcal{P}_{\textup{cubic}}|=38$ different weak Fano threefolds. The reason for this is that the blowup along a curve of type $(0,4)\in\mathcal{P}_{\textup{plane}}\cap\mathcal{P}_{\textup{quadric}}$ gives rise to two different weak Fano threefolds, depending on whether the curve lies on a hyperplane or not. Similarly, the blowup along a curve of type $(3,8)\in\mathcal{P}_{\textup{quadric}}\cap\mathcal{P}_{\textup{cubic}}$ yields two different weak Fano threefolds, depending on whether the curve lies on a hyperquadric section or not. Both of these will become apparent when studying the potential Sarkisov links arising from the weak Fano threefold $X$ in Section 5.
@@ 5. Extremal contractions and Sarkisov links
5.1. Extremal contractions on weak Fano threefolds arising as the blowup of a hyperquadric along a curve
@@ 5.2. Fano threefolds
1, degree 5 and index 1 along a line Cor. 3.27
@@
along a smooth rational curve of degree 4 Cor. 3.27
@@
hyperquadrics in $\mathbb{P}^{4}$ in its singular point Cor. 3.27
@@
0 11 46 $\mathbb{P}^{1}$-bundle over $\mathbb{P}^{2}$ Nr. 31 [ACC+23, Lemma 3.22] Cor. 3.27
@@
22 66 20 conic bundle over $\mathbb{P}^{2}$ Nr. 13 [ACC+23, §5.5] Cor. 3.27
@@ 5.3. Weak Fano threefolds not yielding a Sarkisov link
We have therewith shown that the normal threefold $X^{\prime}$ is indeed only canonical and not terminal. In particular, there exists no Sarkisov link to $X^{\prime}$.
This case can occur for the six pairs $(g,d)\in\{(0,4),\>(2,5),\>(3,8),\>(6,10),\>(8,10),\>(11,12)\}$ (see Table 10). They were all treated in [JPR05] explicitly.
@@
Assume that $C$ does not admit a 4-secant line. Then $C$ is given in $\textup{Pic}(S)$ by $C=5L-E_{1}-E_{2}-E_{3}-E_{4}-3E_{5}$ (4.12+4.13).
@@
Table 10. $(g,d)$-pairs yielding a weak Fano threefold $X$ such that 1) One $K_{X}$-negative extremal ray is the blowup of a smooth hyperquadric $Y\subseteq\mathbb{P}^{4}$ along a smooth irreducible curve $C\subseteq Y$ of genus $g$ and degree $d$. 2) The anticanonical map $\psi$ is a divisorial contraction. More precisely, $\psi$ is the blowup of a non-$\mathbb{Q}$-factorial normal Gorenstein Fano threefold $X^{\prime}$ of index 1 having compound Du Val singularities along a smooth curve $C^{\prime}\subseteq X^{\prime}$ of degree $d^{\prime}$ and genus $g^{\prime}$.
@@ 5.4. Weak Fano threefolds yielding a flop and a Sarkisov link
The construction of each of these weak Fano threefolds $X$ and the Sarkisov links arising from them is given in [JPR11] or [Tak22] for all pairs $(g,d)$, except for the pairs $(g,d)\in\{(0,7),\>(5,9)\}$. For these two pairs, 3.27 guarantees the existence of a corresponding weak Fano threefold and Sarkisov link, which was not known previously.
@@
0 7 10 conic bundle [JPR11, Prop 7.13 Nr. 13] 3.27
@@
5 9 8 conic bundle [JPR11, Prop. 7.13 Nr. 10] 3.27

### v2

@@ 4. Curves on a hyperplane or smooth hyperquadric section
Note that $(0,4)\in\mathcal{P}_{\textup{plane}}\cap\mathcal{P}_{\textup{quadric}}$ since the blowup along a curve of type $(0,4)$ gives rise to two different weak Fano threefolds, depending on whether the curve lies on a hyperplane or not. Similarly, the blowup along a curve of type $(3,8)\in\mathcal{P}_{\textup{quadric}}\cap\mathcal{P}_{\textup{cubic}}$ yields two different weak Fano threefolds, depending on whether the curve lies on a hyperquadric section or not. Both of these will become apparent when studying the potential Sarkisov links arising from the weak Fano threefold $X$ in Section 5.
@@ 5. Extremal contractions and Sarkisov links
5.1. Extremal contractions on weak Fano threefolds of Picard rank two
@@ 5.2. Fano threefolds
1, degree 5 and index 1 along a line Cor. 3.28
@@
along a smooth rational curve of degree 4 Cor. 3.28
@@
hyperquadrics in $\mathbb{P}^{4}$ in its singular point Cor. 3.28
@@
0 11 46 $\mathbb{P}^{1}$-bundle over $\mathbb{P}^{2}$ Nr. 31 [ACC+23, Lemma 3.22] Cor. 3.28
@@
22 66 20 conic bundle over $\mathbb{P}^{2}$ Nr. 13 [ACC+23, §5.5] Cor. 3.28
@@ 5.3. Weak Fano threefolds not yielding a Sarkisov link
We have therewith shown that the normal threefold $X^{\prime}$ is indeed only canonical and not terminal. In particular, there exists no Sarkisov link to $X^{\prime}$. This case can occur for the eight pairs $(g,d)\in\{(0,4),\>(2,5),\>(3,8),\>(4,6),\>(6,10),\>(8,10),\>(11,12),\>(13,12)\}$ (see Table 12). They were all treated in [JPR05] explicitly.
@@
Assume that $C$ does not admit a 4-secant line. Then $C$ is given in $\textup{Pic}(S)$ by $C=5L-E_{1}-E_{2}-E_{3}-E_{4}-3E_{5}$ (4.12 and 4.13).
@@
Note that $C$ admits only finitely many 3-secant-lines since these lines have to be contained in the quartic del Pezzo surface $S$ and there are finitely many lines on $S$ (4.11). To obtain the divisor contracted by $\psi$, we thus really need to consider conics.
Table 11. $(g,d)$-pairs yielding a weak Fano threefold $X$ such that 1) One $K_{X}$-negative extremal ray is the blowup of a smooth hyperquadric $Y\subseteq\mathbb{P}^{4}$ along a smooth irreducible curve $C\subseteq Y$ of genus $g$ and degree $d$. 2) The anticanonical map $\psi:X\rightarrow X^{\prime}$ is a divisorial contraction to a smooth curve. More precisely, $\psi$ is the blowup of a non-$\mathbb{Q}$-factorial normal Gorenstein Fano threefold $X^{\prime}$ of index 1 having compound Du Val singularities along a smooth curve $C^{\prime}\subseteq X^{\prime}$ of degree $d^{\prime}$ and genus $g^{\prime}$.
@@
Table 12. $(g,d)$-pairs yielding a weak Fano threefold $X$ such that 1) One $K_{X}$-negative extremal ray is the blowup of a smooth hyperquadric $Y\subseteq\mathbb{P}^{4}$ along a smooth irreducible curve $C\subseteq Y$ of genus $g$ and degree $d$. 2) The anticanonical map $\psi:X\rightarrow X^{\prime}$ is a divisorial contraction to a point. More precisely, $\psi$ is the blowup of a non-$\mathbb{Q}$-factorial canonical Gorenstein Fano threefold $X^{\prime}$ from the list [JPR05, A.1] in a point.
$g$ $d$ $K_{X}^{3}$ Reference + Existence
4 6 24 [JPR05, A.4 Nr. 25] with $r=3,H^{3}=2$ and $r^{\prime}=1$
13 12 6 [JPR05, A.4 Nr. 25] with $r=3,H^{3}=2$ and $r^{\prime}=2$
@@ 5.4. Weak Fano threefolds yielding a flop and a Sarkisov link
The construction of each of these weak Fano threefolds $X$ and the Sarkisov links arising from them is given in [JPR11] or [Tak22] for all pairs $(g,d)$, except for the pairs $(g,d)\in\{(0,7),\>(5,9)\}$. For these two pairs, 3.28 guarantees the existence of a corresponding weak Fano threefold and Sarkisov link, which was not known previously.
@@
0 7 10 conic bundle [JPR11, Prop 7.13 Nr. 13] 3.28
@@
5 9 8 conic bundle [JPR11, Prop. 7.13 Nr. 10] 3.28